OpenStudy (anonymous):

the difference of two numbers is 19.the larger number is 3more than twice the smaller.find the numbers

OpenStudy (anonymous):

x-y=19 x=2y+3 substituting x=2y+3 in second equation 2y+3-y=19 y=16 x-16=19 x=35 numbers are 16 and 35